OpenStudy (anonymous):

determine the number of fl oz in a 2 liter container

OpenStudy (abcdefghijklmq):

about 67.628

OpenStudy (anonymous):

how did you get that?

OpenStudy (zpupster):

\[(\frac{ 2L }{ ? })(\frac{ 1qt }{ 1.06L})(\frac{ 32oz }{ 1qt. })\]